You can put the tags on the left with the legno option.

You can align equations with automatic numbering each line.

(1)  2x + 3 = 5
(2)  2x + 3 − 3 = 5 − 3
(3)  2x = 2
(4)  \frac{2x}{2} = \frac{2}{2}
(5)  x = 1

You can suppress numbering.

2x + 3 = 5
2x + 3 − 3 = 5 − 3
2x = 2
\frac{2x}{2} = \frac{2}{2}
x = 1

You can have subnumbering.

(6a)  2x + 3 = 5
(6b)  2x + 3 − 3 = 5 − 3
(6c)  2x = 2
(6d)  \frac{2x}{2} = \frac{2}{2}
(6e)  x = 1

You can interject text.

2x + 3 = 5
2x + 3 − 3 = 5 − 3  Subract 3 from both sides
2x = 2  Simplify
\frac{2x}{2} = \frac{2}{2}  Divide both sides by 2
x = 1  Simplify

You can align more than one set of equations.

2x + 3 = 5  r = 0
2x + 3 − 3 = 5 − 3  1 + 2 sin θ = 0
2x = 2  sin θ = −\frac{1}{2}
\frac{2x}{2} = \frac{2}{2}  θ = \frac{7π}{6}
x = 1